### v1

# Thai Winograd Schemas: A Benchmark for Thai Commonsense Reasoning

Abstract
Commonsense reasoning is one of the important aspect of natural language understanding, with several benchmarks developed to evaluate it. However, only a few of these benchmarks are available in languages other than English. Developing parallel benchmarks facilitates cross-lingual evaluation, enabling a better understanding of different languages. This research introduces a collection of Winograd Schemas in Thai, a novel dataset designed to evaluate commonsense reasoning capabilities in the context of the Thai language.
Through a methodology involving native speakers, professional translators, and thorough validation, the schemas aim to closely reflect Thai language nuances, idioms, and cultural references while maintaining ambiguity and commonsense challenges. We evaluate the performance of popular large language models on this benchmark, revealing their strengths, limitations, and providing insights into the current state-of-the-art. Results indicate that while models like GPT-4 and Claude-3-Opus achieve high accuracy in English, their performance significantly drops in Thai, highlighting the need for further advancements in multilingual commonsense reasoning.
Thai Winograd Schemas: A Benchmark for Thai Commonsense Reasoning
Phakphum Artkaew Department of Electrical and Computer Engineering Tandon School of Engineering New York University pa2497@nyu.edu
1 Introduction
Commonsense reasoning is an important challenge in artificial intelligence; however, most resources and benchmarks are in English, with only a few translations Davis (2023). The Winograd Schemas Challenge (WSC) has emerged as a widely adopted benchmark for evaluating the commonsense reasoning capabilities of language models Levesque et al. (2012). WSC is featured in both GLUE and SuperGLUE. These benchmarks are widely used to evaluate a model’s understanding of general language Wang et al. (2018) Wang et al. (2019). Significant progress has been made in developing commonsense reasoning models for high-resource languages like English. However, their performance on low-resource languages remains largely unexplored. This research gap hinders the development and fair evaluation of multilingual and cross-lingual NLU systems. In their work, Linzen (2020) notes, “We should ensure that our architectures, like human learners, are not optimized for English Bender (2011). To do so, we should develop matched training corpora and benchmarks for multiple languages. A composite score could reflect average performance across languages Hu et al. (2020).” Consequently, the development of such benchmarks or corpora is vital for tasks evaluating the current state of systems across languages.
There have been several translations of the Winograd Schemas into various languages, including French Amsili and Seminck (2017), Portuguese De Melo et al. (2019), Mandarin Bernard and Han (2020), Hebrew Shwartz (2024), Hungarian Vadász and Ligeti-Nagy (2022), and Russian Shavrina et al. (2020). Before our translation, no Thai version of the Winograd Schemas existed. This lack prompted us to take the initiative.
The key contributions of our work are:
1.
We introduce Thai-WS, the first benchmark dataset for evaluating the Winograd task and commonsense reasoning capabilities of language models and NLP systems in the Thai language.
2.
We evaluated the performance of state-of-the-art models, including GPT, Claude, and Typhoon, on the Thai-WS dataset to assess their cross-lingual reasoning abilities.
2 Related Work
Background
The Winograd Schema Challenge, introduced by Levesque et al. (2012) in 2012 as an alternative Turing Test, consists of pairs of similar sentences with slight differences that introduce ambiguities resolved through general knowledge and logical reasoning. A classic example by Winograd (1972) illustrates the importance of context: "The city councilmen refused the demonstrators a permit because they feared/advocated violence." The choice of "feared" or "advocated" shifts the referent of "they" from the city councilmen to the demonstrators. This is because "feared" aligns with concerns typical of councilmen’s reasons for refusal, whereas "advocated" suggests demonstrators’ motivations, reflecting how context dictates meaning.
Other Thai Reasoning Benchmarks
To the best of our knowledge, three public benchmarks are available for Thai reasoning. First is XNLI, which is a translation of the MNLI benchmark into Thai, among 14 other languages Conneau et al. (2018). Second is XCOPA, translated from the COPA benchmark to assess causal commonsense reasoning Ponti et al. (2020). Third, the recently released M3Exam, is based on examination questions Zhang et al. (2024). Unlike these benchmarks, the Winograd Schema Challenge specifically assesses commonsense reasoning through pronoun reference disambiguation task.
A Simple Method for Commonsense Reasoning
Trinh and Le (2018) introduced a simple approach to solve the Winograd Schema hallenge by substituting one of the pronouns in the sentence and allowing the language model to determine which substitution has a higher probability. This method effectively frames Winograd Schemas as a binary classification task. Since then, many language models, such as GPT-2 Radford et al. (2019) and GPT-3 Brown et al. (2020), have used this method to evaluate their performance on Winograd Schemas. However, due to limited access to models and computational power, the evaluation of this version of the Winograd Schema Challenge was conducted differently, using a prompt-based method, which will be discussed in the experimental setup section.
3 Dataset Construction
3.1 Translation
Two professional translators, who were native Thai speakers fluent in English and had experience translating from English to Thai, were hired. In a pilot translation phase, one native speaker translated the first 85 sentences. Based on a qualitative analysis of these 85 sentences, guidelines were provided for a second native speaker to translate the remaining 200 sentences. In total, 285 sentences were translated from English to Thai. Translation guidelines were provided, instructing them to adapt names and contexts to suit the Thai language while preserving the ambiguity and nuances of the original schema. The translators were also asked to mark any translated names and translations they were unsure about, so that the validator in the next step could pay extra attention to those instances. For example, in Figure 1, the names Paul and George were changed to Mana and Piti, respectively, adapting the names to better suit the Thai context while preserving the essence of the original content.
Some phrases could be directly translated, while others required adjustments to names and contextual adaptations to better suit the Thai language and culture. However, there were a few instances where the translators highlighted certain phrases in red, indicating that they are worth mentioning.
For example:
Many people start to read Paul’s books and can’t put them down. They are gripped because Paul writes so well.
Many people start to read Paul’s books and can’t put them down. They are popular because Paul writes so well.
Fred covered his eyes with his hands, because the wind was blowing sand around. He opened them when the wind stopped.
Fred covered his eyes with his hands, because the wind was blowing sand around. He lowered them when the wind stopped.
Lily spoke to Donna, breaking her silence.
Lily spoke to Donna, breaking her concentration.
Most of the translation problems encountered involve examples like 1, where in English the word “they” can refer to both people and objects. In this case, “they” can mean either “people” or “Paul’s books.” However, in Thai, it is uncommon to use the same pronoun for both objects and people. Therefore, an alternative pronoun that is acceptable in Thai was chosen. In this instance, the English equivalent word “those” was used instead.
Similarly, in 2, “them” can refer to either “his eyes” or “his hands.” In Thai, it is not common to use a pronoun in this way. The translation was changed to the English equivalent of “those two,” which sounds more natural in Thai.
In other cases, the meaning had to be adjusted to sound natural in Thai. For example, 3 “breaking her silence” is not a common expression in Thai, so it was revised to “The silence was replaced by her voice.”
3.2 Validation
| Model | Accuracy (English) | Accuracy (Thai) |
| Typhoon | 64.56% | 54.04% |
| Claude-3-Haiku | 62.81% | 51.58% |
| Claude-3-Sonnet | 80.70% | 63.51% |
| Claude-3-Opus | 92.63% | 77.54% |
| GPT-3.5 | 71.93% | 51.23% |
| GPT-4 | 93.68% | 72.28% |
| Human | 90% | -% |
The translated Winograd Schemas were reviewed with three native Thai speakers, and final adjustments were made to ensure clarity. A validator was provided with the translations and tasked with identifying any potential issues. They were instructed to pay closer attention to the text marked by the translators, which included changes to names and translations that the translators were unsure about. Based on the validator’s feedback, final adjustments to the translations were made, and typographical errors were corrected. Furthermore, the dataset will be made publicly available111Avaliable at https://huggingface.co/datasets/pakphum/winograd_th., inviting other native Thai speakers to verify it and suggest any necessary adjustments.
4 Experimental Setup
Large language models were evaluated on both Thai and English Winograd Schemas to assess Thai natural language processing and enable cross-linguistic comparison of commonsense reasoning. It is important to note that there are two versions of the Winograd Schema Challenge (WSC): WSC-273 and WSC-285. In this study, WSC285 was used.
Typhoon
The Thai large language model, named Typhoon, is built on the Mistral-7B architecture Pipatanakul et al. (2023). Mistral-7B features grouped-query attention for faster responses and sliding window attention for cost-effective processing of any text length Jiang et al. (2023). Typhoon was pre-trained on the MC4 and OSCAR datasets, which include approximately 3 TB of Thai text. A Thai subword tokenizer was added to enhance its performance on Thai language tasks. In this experiment, the specific model tested was typhoon-instruct-0219.
Claude
Anthropic’s large language model Anthropic (2024a), designed to be a helpful and honest assistant, has its model architecture details kept private. In this experiment, all the Claude-3 models were tested to compare performance. The specific models tested were claude-3-haiku-20240307, claude-3-sonnet-20240229, and claude-3-opus-20240229.
GPT
GPT or Generative Pre-trained is one of the most popular and powerful large language models available. While the specific details of GPT-4’s architecture are not publicly disclosed, it can be assumed that it represents an advancement over GPT-3. GPT-3 is an autoregressive, decoder-only model with 175 billion parameters Brown et al. (2020). Although the technical report does not specify its performance in Thai specifically, it asserts that GPT-4 surpasses GPT-3 in the MMLU task for the Thai language Achiam et al. (2023). In this experiment, gpt-4-0613 and gpt-3.5-turbo-0125 were tested.
Implementation and evaluation
Language models were evaluated using a prompt structure approach. A system prompt and user prompt structure were utilized to assess the models. A system prompt provides context and instructions to language models before a task, specifying the model’s role, personality, tone, or other relevant information to enhance its responses Anthropic (2024b). The system prompt was designed to prepare the model for the Winograd Schemas task, instructing it to respond with the correct answer. The evaluation focused on accuracy, counting only answers that exactly matched the correct answer in the corresponding schemas. Outputs such as “the answer is the city councilmen ” were not counted as correct; only precise answers like “the city councilmen" were considered correct. The models were evaluated through their respective APIs. More details on the evaluation and prompt setup can be found in A.
5 Result and Discussion
Table 1 shows the accuracy results when performed on each model. There were several human baseline for the English Winograd Kocijan et al. (2023). Our English human baseline is derived from the study conducted by Davis et al. (2016), which involved human participants evaluating Winograd Schemas in English. In Davis et al. (2016) research, human accuracy was observed to be approximately 90%. The human baseline for Thai has not been conducted yet.
Comparison of models
Typhoon achieved an accuracy of 64.56% in English and 54.04% in Thai, showing a drop in performance when handling Thai schemas. Claude-3-Haiku scored 62.81% in English and 51.58% in Thai, similarly reflecting a lower accuracy in the Thai language. Claude-3-Sonnet demonstrated better performance with 80.70% accuracy in English and 64.56% in Thai. Claude-3-Opus achieved high accuracy of 92.63% in English and 77.54% in Thai, showing the best performance among the Claude-3 variants.
GPT-4 scored the highest in English with 93.33% and performed well in Thai with 73.68% accuracy. GPT-3.5 showed a significant drop from GPT-4, with 71.58% accuracy in English and 51.58% in Thai.
Overall, the results indicate that all models perform better on the English version of the Winograd Schema Challenge compared to the Thai version. The performance drop in Thai suggests that these models may have more difficulty handling the nuances and linguistic structures of Thai.
6 Conclusion
In conclusion, the Thai Winograd Schemas benchmark represents a noteworthy step in evaluating commonsense reasoning capabilities within the Thai language context. This novel dataset, meticulously developed and validated by native speakers and professional translators, aims to preserve linguistic and cultural nuances unique to Thai. The comprehensive evaluation of state-of-the-art language models, including GPT-4, Claude-3 variants, and Typhoon, on both English and Thai versions of the Winograd Schema Challenge offers insights into their cross-lingual performance. The observed performance drop in Thai suggests challenges these models may face in handling low-resource languages, indicating a need for further research and development in multilingual natural language processing.
Limitations
While cultural nuances are preserved as much as possible during the translation process, it is acknowledged that complete preservation is not always achievable due to differences between Thai and English. Additionally, there is a limitation regarding the Thai human baseline for Thai Winograd Schemas, as such a baseline has not yet been established. Direct evaluation on large language models like GPT-4 and Claude-3 cannot be performed due to lack of access and insufficient computational power. Therefore, an alternative approach using prompt-based evaluation is adopted.
Acknowledgments
Thank you to the two translators, Chanikarn Inthongpan and Korakoch Rienmek, who worked on translating these Winograd schemas. It was challenging to find individuals skilled enough for this specific task, which demands a deep understanding of language. I also extend my gratitude to my friends and professors who supported me throughout the writing process.
Appendix A Prompt Evaluation
To ensure consistency and reproducibility in evaluating each model, specific settings were implemented alongside the prompt structure. The system prompt was devised to provide clear instructions: “You will be provided with a sentence and a snippet containing a pronoun enclosed in asterisks (**). Your task is to determine the correct referent of the pronoun from the given options. Respond only with one of the provided choices, exactly as it is written. For example, if the options are ‘The city councilmen’ and ‘The demonstrators’, respond only with ‘The city councilmen’ or ‘The demonstrators’.” Additionally, the temperature parameter was set to 0 to minimize randomness, and when possible, the seed value was fixed at 77 to further enhance reproducibility. For the user’s prompt, the model was given a Winograd Schema Challenge, formatted with the scenario and its options separated by a new line, adhering to the same standards for consistency and replicability.
[…]
Only exact matches to the correct choice were counted as correct answers. For instance, responses such as “the answer is The city councilmen” were not counted as correct; only precise answers like “The city councilmen” were considered correct. This process was executed through the model API, ensuring that each evaluation followed the same method and settings for all models.
The code used to reproduce all of our experimental results is available at the following GitHub repository: https://github.com/PhakphumAdev/Thai-Winograd.